There are 12 female students and 18 male students in a class. Every student independently has a probability 0.3 of being a Math major. There are 14 Math majors in the class. Find the probability that among the Math majors, 6 students are female.
Answer #1

since probability of being a math major is independent of gender

P( 6 students are female )=(choose 6 from 12 female and 8 from 18 male)

=(12C6)*(18C8)/(30C14) =924*43758/145422675 =0.2780